OpenStudy (anonymous):

radical 44 minus radical 11? O.o

OpenStudy (anonymous):

not sure what you mean...drawing it might help?

OpenStudy (anonymous):

\[\sqrt{44}- \sqrt{11}\]

OpenStudy (saifoo.khan):

\[\sqrt{44} \to \sqrt{4 \times 11} \to 2 \sqrt{11}\]

OpenStudy (anonymous):

wait a minute I am doing operations with radical expressions so is this right?

OpenStudy (saifoo.khan):

Yes. that isn't the final answer though.

OpenStudy (anonymous):

how do you get the final answer?

OpenStudy (saifoo.khan):

Now we solved the first term, we are left with; \[2 \sqrt{11} - 1\sqrt{11}\]

OpenStudy (anonymous):

oh ok so the answer is just:\[\sqrt{11}\]

OpenStudy (saifoo.khan):

Yes sir.

OpenStudy (anonymous):

ma'am.

OpenStudy (saifoo.khan):

Ops. ma'am*. ;)